OpenStudy (anonymous):

Jill is 6 years older than her brother ryan.write an equation relating jills age to ryans age . explain what each varible represents.use your equation find ryans age when jill is 14.

OpenStudy (mertsj):

j=jill's age r = ryan's age j=r+6 14=r+6 8=r